# Pangenome-guided sequence assembly via binary optimisation

### v1

@@ 3 Discussion
In this work, we initiate a study on using pangenomes to aid in the assembly of short-read data. We adopt an optimisation approach to tackle assembly, which we believe is well-suited to mitigate the noise inherent in short-read sampling. We demonstrate that our methods enhance the accuracy of the assembly while significantly reducing the number of contigs reported compared to state-of-the-art de novo assemblers.
@@
Our pipeline consists of 2 main stages: graph annotation and path finding. We investigate the impact on solution quality of using different software for both steps.
@@
For the graph annotation, we find that minigraph generally performs strongly, minimising the number of breaks in the assembly as well as the number of large indels, at the cost of reporting several contigs per assembly. Usually, we prefer several accurate contigs over fewer, more inaccurate ones.
@@
For the solvers, we see that our (classical) binary optimisation approaches are at the very least competitive with the bleeding-edge bioinformatics tool pathfinder. As the problem size increases and the exhaustive search of pathfinder is no longer feasible, optimisation methods may be better placed to report reasonable assemblies.
@@
We also investigate the feasibility of quantum computing in solving these problems. We first test a QA approach on a set of middling-complexity problems, observing that the solution quality is broadly similar to pathfinder, even though only short time limits were used. This is a positive sign for the potential of using QA to solve industry-scale problems in the future, as the hardware continues to improve. Finally, we simulate the QAOA on some small problems as a first indicator of the potential for standard, circuit model quantum computers to solve pangenomic problems. We are able to converge a QAOA experiment at this small scale. While performing a hardware experiment at scale is required to make any serious claims about quantum utility in this domain, we see this as a promising proof-of-principle.
One drawback of the current optimisation formulation is the lack of edge information. When we have graphs with multiple paths through the same set of nodes, it can be impossible to distinguish nodes A+ B+ C+ from A+ B- C+. Pathfinder combines edge weights with node weights to resolve these conflicts.
While it is straightforward to include an edge weight term in the abstract optimisation cost functions of eqs. 2, 3 and 4, it is not clear how to translate such terms into a QUBO. The information of whether an edge was traversed at a certain time is inherently quadratic, given by terms $x_{t,\,i}x_{t+1,\,j}$. To obtain a cost function that is minimised when the number of traversals of an edge is close to the weight of that edge requires terms like $\sum_{t}\big{(}x_{t,\,i}x_{t+1,\,j}-w(i,\,j)\big{)}^{2}$ for each $(i,\,j)\in E$. These higher-order terms incur overheads in the number of variables if they are re-expressed as quadratic terms. A local graph preprocessing step to remove zero-weight edges that do not break graph connectivity is likely to achieve an approximate equivalent without the need to incorporate edge data directly into the QUBO model.
All graph path finding methods are impacted by the quality of the input kmer counts. Further improvements can be made to these tools, particularly concerning non-unique mapping, which would improve the quality of all solvers. However, there is further missing information yet to be considered where larger sequence fragments trace a path through many small nodes, forming longer-range constraints rather than the nearest neighbour ones encoded in the graph edge counts.
The kmer mapping software used to generate node weights does not use Illumina read-pairing metrics. In a traditional linear reference alignment, the read pairing is an important factor for resolving some repeats, and the same could also be applied for pangenome mapping. Pairing information could potentially also be used in inferring paths, utilising knowledge of the template size distribution.
One possible direction that improves resource usage and allows edge weight data to be included is Higher-order Unconstrained Binary Optimisation (HUBO) variants. HUBO formulations are more expressive than QUBO, with terms containing products of arbitrarily many binary variables; the downside is that (classically) solving them is generally more difficult and less well-explored in the literature. Notably, however, HUBO may not have such a solution overhead when mapped to QAOA algorithms on quantum computers. In that setting, each layer of the cost Hamiltonian now contains multi-qubit $Z$ rotations rather than single- and two-qubit rotations only. These gates need to be compiled down to 2-qubit gates, but constructions that are linear in the number of qubits involved exist [22]. We could also consider Prog-QAOA [23] formulations, which directly encode optimisation problems into QAOA cost Hamiltonians, without going through a QUBO or HUBO step. We expect that either of these 2 encodings could achieve $\mathcal{O}(N\log(N))$ qubit scaling, at the cost of increased circuit depth.
$C_{G}(W)=\sum_{v\in V}\log(L(v))\Bigl{(}\#W(v)-w(v)\Bigr{)}^{2},$ (1)
where $L(v)$ is the length of the sequence stored in node $v$.
@@ 4.2 Pangenome creation
There is more than one type of pangenome graph that could be constructed.
A de Bruijn graph approach would make every exact repeat into a crunch point, with a single node having many edges in and out. Matching kmers against a de Bruijn graph is easy, as each kmer is likely only present in a single graph node (assuming the match kmer length and graph-construction kmer length are similar). Tracking edge transitions between consecutive kmers may still provide a single route through this graph, but it is a complex task and requires knowledge of the history of edges (path fragments).
Tools such as minigraph prefer to build a pangenome that is broadly co-linear with the input genomes. This reduces the number of tangles and loops in the graph, but it means repeated sequence may end up in multiple nodes. This is the style of pangenome published by the HPRC, and it has the benefit of being able to thread standard linear reference coordinates, such as GRCh38, through the pangenome graph. Therefore, this is the style of pangenome graph we focus on.
A pangenome can also be complete or sparse: it could include every single base difference used in the input data, which leads to a large graph with many nodes and edges, or it could collapse small localised variations into a single node. In this case, the node sequence will be either a representative single sequence or a consensus of all the donor genomes that passed through that node during graph construction. Minigraph takes this latter approach, and it is important for us to keep the graph complexity as low as we can for the problem size to be amenable to existing quantum computers.
We build a simulated genome population using our own tool genome_create with fixed random seeds for reproducibility. The initial genome contains Short-Tandem Repeats (STRs), longer Copy Number Variations (CNVs), short and long repeat elements, translocations and inversions, and random point mutations. Each subsequent population member is haploid, being derived from a single randomly chosen previous genome. This is then further modified by the same types of variation listed above, but at a substantially reduced rate, giving rise to a population of similar sequences with a common complex structure that can form population graphs. We use a subset of our population to build the pangenome graph using minigraph. The remaining sequences are used to simulate shotgun sequencing by replicating to 30-fold coverage and fragmenting into short sequences, with a uniform distribution of sequencing errors.
@@ 4.3 Annotating the graph with copy numbers
For each of the remaining members of the simulated population, we produce fake short-read sequencing data. For simplicity, this is single-ended (i.e. not in pairs from a larger template). They are randomly distributed, from either strand, and with a random distribution of base substitution errors. No indel errors are modelled.
There are existing tools that can take a short sequence, align it against a graph (GFA) and record the path it took through the nodes. These include GraphAligner, vg giraffe and minigraph. We also have our own kmer2node tool. All four tools annotate the GFA with edge traversal counts and kmer counts within node sequences. The presence of repeats can still be problematic, as there may be multiple possible paths. It may not be possible to distinguish between genuine sequence rearrangements and secondary hits caused by multiple copies of the repeated sequence in the pangenome. When reported, we filter out secondary alignments.
As mentioned above, our pangenome graph node sequences are a single representative or consensus sequence. However, the training genomes used to construct that pangenome may have many point mutations and small indels, which reduce kmer match rates. Kmer2node may use a list of recognition sequences instead of a single one, so we use GraphAligner to get end-to-end alignment of the graph training data to produce a list of recognition sequences per node instead of the recorded single one. This improves our kmer hit rate, but may also decrease the percentage that uniquely map. While indexing the GFA, we also get an estimate of the percentage of kmers we expect to be unique. This provides us with a kmer hit-rate expectation during short-read alignment of short-read data, allowing us to improve the assessment of node sequence depth.
All the tools produce a GFA with kmer counts per node and frequencies of edge traversals. For pathfinder, this is sufficient as it can convert these to depth-normalised counts. To aid consistency, we also use pathfinder’s sub-graph connection analysis and depth-normalised counts as input to the QUBO solvers, although note this does not utilise edge count information.
@@ 4.4 Classical post-processing: evaluating path solutions
Our path solvers produce a list of nodes and their orientation. From this, we can concatenate the sequences recorded in the GFA nodes, reverse-complementing as necessary, to produce a candidate genome sequence. Note that even if the path is perfect, this candidate genome may not be identical to the true genome, as our GFA files produced by minigraph collapse small variations into a single node instead of producing a much larger graph with every point mutation present. It is also possible that the true genome had novel structural rearrangements (copy-number changes or translocations) that were not present in any of the training data used to produce the pangenome graph, meaning a perfect path does not exist.
We align this candidate genome sequence against the known truth using bwa mem, reporting only primary and supplementary hits. This produces one or more alignments, each representing co-linear matches between the candidate and true genomes, being broadly equivalent to the contigs produced by a de novo assembler. Each alignment record may, in turn, have smaller variations (substitutions and small indels) that were not deemed large enough to break the alignment into two. This gives us a unified way to compare pangenome path reconstruction against more traditional de novo sequence assembly tools. We report the number of contigs (incorrect path forks or large indels causing a supplementary alignment), small indels (>=>= 10bp), and regions of substantial base variation (multi-nucleotide polymorphism of 30% variance within a 100bp window). Additionally, we count the percentage of bases in the true genome that are covered by the candidate genome and vice versa, so we can assess coverage statistics.
A further optional refinement step is to realign the short reads back to the generated candidate genome sequence and to take the consensus of the newly aligned sequences. This corrects for novel mutations in this sample or simplifications made during the minigraph pangenome creation step.
@@ 4.5 Formulating sequence assembly as an optimisation problem
After annotation, our data is a graph with non-negative weights on the nodes and edges. Our current formulation only uses node weights, whereas tools such as pathfinder are able to use both sets of weights. Extending our methodology to include edge weights is a future direction that we discuss in section 3.
We formally define the pangenome graph annotated with copy numbers as a triplet $G=(V,E,w)$, where $V$ is the set of vertices, $E\subseteq V\times V$ is the set of edges and $w:V\rightarrow\mathbb{R}_{+}$ is a weight function giving the copy number of each vertex.
We then seek a walk on the graph that best explains the copy number data. This graph-traversal problems shares superficial similarities with the Travelling Salesman Problem. In particular, we define the following optimisation problem, which we refer to as Tangle Resolution due to the often complex, knot-like structures present in pangenomes. Note the implicit requirement in the problem that each step in $W$ traverses an edge in $G$.
@@
While this model serves as a useful starting point, it does not encompass all the biological features. Most notably, it neglects to take into account the double-stranded nature of genomic data, wherein both a sequence and its reverse complement are present in the DNA. To better reflect the biology, we must consider a directed graph. The vertex set is now formed of 2 halves $V_{+},V_{-}$ The sequence data corresponding to $v_{-}\in V_{-}$ is then the reverse complement of the sequence data corresponding to $v_{+}\in V_{+}$. Edges of the graph are directed and come in pairs: an edge $A+\rightarrow B+$ implies the existence of an edge $B-\rightarrow A$-, and vice versa. We call such a graph $G=(V_{+}\cup V_{-},E,w)$ oriented.
@@
Finally, for the setting of human data, we need to resolve diploid DNA. We therefore need to find a pair of walks $W_{1},\,W_{2}$ through an oriented graph that combine to explain the data. We give a formal problem statement of Diploid Tangle Resolution:
@@ 4.6 Mapping guided alignment to binary optimisation
The QUBO formulation defines a class of optimisation problems characterised by the minimisation of a quadratic polynomial over binary variables $x_{i}$}. Since $x_{i}^{2}=x_{i}$ for $x_{i}\in\{0,1\}$, any linear term can be replaced by a quadratic one. Also, any constant offset can be neglected since it just shifts the energy landscape without affecting the optimal assignment. Therefore, the objective can be formally defined as minimising $C(\{x_{i}\})=x^{T}Mx$, where $x\in\{0,1\}^{n}$ and $M\in\mathbb{R}^{n\times n}$ is a symmetric, real-valued square matrix. QUBO problems enjoy a high degree of expressivity, being able to represent classic combinatorial optimisation problems such as MAX-CUT and the Travelling Salesman Problem (TSP), while also remaining tractable to modern solvers at moderate sizes. Solution techniques include generic branch-and-bound methods, as employed by commercial solvers such as Gurobi, as well as domain-specific heuristics used by the open source library MQLib.
QUBO is equivalent, under a linear change of variables, to the Ising model in statistical physics, a formulation extensively used in quantum annealing devices such as those developed by D-Wave Systems. As such, QUBO is widely employed in quantum computing contexts, where its unconstrained nature and binary domain align with the hardware capabilities of current quantum and neuromorphic processors.
The primary advantages of QUBO include its flexibility in encoding diverse problems within a unified mathematical framework and its compatibility with both quantum and classical heuristic solvers. Its unconstrained nature simplifies solver implementation by eliminating the need for explicit constraint handling, instead requiring constraint satisfaction to be incorporated into the objective function through penalty terms. Notably, however, the QUBO forms of many classic graph-traversal problems, such as the Travelling Salesman Problem, require quadratically many variables in the graph size, with each variable encoding whether the path visits a certain node at a certain time. QUBO formulations of these tasks traditionally use binary variables $x$ with 2 indices $t$ and $v$, with the variable $x_{t,v}=1$ if the solution path visits vertex $v$ at time $t$. The cost function is easily formulated by summing the cost of each step $c_{vv^{\prime}}$ over the edges that are traversed, leading to $\sum_{t,v,v^{\prime}}c_{vv^{\prime}}x_{t,v}x_{t+1,v^{\prime}}$ For an assignment to represent a valid path, it is necessary to impose the constraint that exactly one vertex is visited at each time. This is accomplished through adding Lagrange multiplier terms to the cost function for each $t$, taking the form $\Bigl{(}\sum_{v\in V}x_{t,v}-1\Bigr{)}^{2}$.
@@
The total cost function consists of 3 parts: the first is a Lagrange multiplier term to enforce that the variable assignment encodes a walk that visits exactly 1 vertex at each time; the second is a multiplier term to enforce that, at each step the walk traverses a graph edge, including travelling to or remaining in the virtual “end” node; the final term encodes the cost function $C_{G}(W)$.
@@
$C^{2}_{G}(\{x_{t,v}\})$ $=\Lambda_{2}\sum_{t=1}^{T-1}\left(\sum_{v,v^{\prime}\in V}x_{t,v}x_{t+1,v^{\prime}}\mathbbm{1}_{\{(v,v^{\prime})\notin E\}}+\sum_{v\in V}x_{t,\text{end}}x_{t+1,v}\right),$ (6)
@@
There is some art to choosing the values of the Lagrange multipliers $\Lambda_{1},\Lambda_{2}$. If they are sufficiently large, then we can guarantee that the assignment that minimises $C_{G}(\{x_{t,v}\})$ encodes a walk that minimises $C_{G}(W)$. However, if they are too large, then the energy landscape will have steep valleys around assignments that satisfy the constraints, and changes in the cost function will be nearly indistinguishable to the optimiser. We find that intermediate-size values of $\Lambda_{1}=10$ and $\Lambda_{2}=5$ work well in practice.
To extend this method to Oriented Tangle Resolution, we simply introduce a pair of variables for each vertex and each time, resulting in $(2N+1)T$ variables. For Diploid Tangle Resolution, we further introduce a new set of variables for the second path, resulting in $2(2N+1)T$ variables.

### v2

@@ 3 Discussion
In this work, we initiate a study on using pangenomes to aid in the assembly of short-read data. We adopt a novel optimisation approach to tackle assembly, which we believe is well-suited to mitigate the noise inherent in short-read sampling. We demonstrate that our methods enhance the accuracy of the assembly while significantly reducing the number of contigs reported compared to state-of-the-art de novo assemblers. The optimisation is designed to be amenable to both classical and quantum computing approaches and scales more favourably with problem size than existing methods.
Moreover, in challenging cases such as repetitive regions, our approach is more robust since errors during the copy number estimation step will be less impactful on our cost function approach compared to the existing state-of-the-art pathfinder, whose copy number assignment is prescriptive. For such instances, all sequence assembly methods will suffer from decreased accuracy due to the difficulty of assigning kmer counts; our kmer2node tool includes mitigation strategies, which we discuss in section S3.2.
@@
Our pipeline consists of 3 main stages: graph annotation, copy number estimation and path finding. We investigate the impact on solution quality of using different software for both the graph annotation and path finding steps. For copy number estimation, we only use the capabilities of pathfinder. Further exploration in this direction could certainly be fruitful; one option might include training a machine-learning model to perform this task.
@@
For the graph annotation, we find that minigraph generally performs strongly, minimising the number of breaks in the assembly as well as the number of large indels, at the cost of reporting several contigs per assembly. We usually prefer several accurate contigs over fewer, more inaccurate ones.
@@
For the solvers, we see that our (classical) binary optimisation approaches are at the very least competitive with the bleeding-edge bioinformatics tool pathfinder. On the intermediate-size instances we test, the classical optimisers report good results within reasonably short time limits, often reporting better solutions than pathfinder on specific problems. They could be included as a parallel assembly method when high-quality results are desired, taking a consensus of all the candidate solutions as the final answer. As the problem size increases and the exhaustive search of pathfinder is no longer feasible, optimisation methods may be better placed to report reasonable assemblies.
@@
We also investigate the feasibility of quantum computing in solving these problems. We first test a QA approach on a set of middling-complexity problems, observing that the solution quality is broadly similar to pathfinder, even though only short time limits were used. This is a positive sign for the potential of using QA to solve industry-scale problems in the future, as hardware continues to improve. However, there is currently no evidence of utility in using these systems for our problem.
Finally, we simulate the QAOA on some small problems as an initial indicator of the potential for standard, circuit-model quantum computers to solve pangenomic problems. We can converge a QAOA experiment at this small size. While performing a hardware experiment at industry scale is required to make any serious claims about quantum utility in this domain, we see this as a promising proof-of-principle.
For a discussion of possible future research directions, see section S2.
@@ 4.2 Pangenome creation
Tools such as minigraph prefer to build a pangenome that is broadly co-linear with the input genomes. This reduces the number of tangles and loops in the graph, but it means repeated sequence may end up in multiple nodes. This is the style of pangenome published by the HPRC, which has the benefit of allowing standard linear reference coordinates, such as GRCh38, to be threaded through the pangenome graph. Therefore, this is the style of pangenome graph we focus on.
A pangenome can also be complete or sparse: it could include every single base difference used in the input data, which leads to a large graph with many nodes and edges, or it could collapse small localised variations into a single node. In this case, the node sequence will be either a representative single sequence or a consensus of all the donor genomes that passed through that node during graph construction. Minigraph takes this latter approach, which is useful since we need to keep the graph complexity low for the problem size to be amenable to existing quantum computers.
We build a simulated genome population using our own tool genome_create with fixed random seeds for reproducibility. The initial genome contains Short-Tandem Repeats, longer CNVs, short and long repeat elements, translocations and inversions, and random point mutations. Each subsequent population member is haploid, being derived from a single randomly chosen previous genome. This is then further modified by the same types of variation listed above, but at a substantially reduced rate, giving rise to a population of similar sequences with a common complex structure that can form population graphs. We use a subset of our population to build the pangenome graph using minigraph. The remaining sequences are used to simulate shotgun sequencing by replicating to 30-fold coverage and fragmenting into short sequences, with a uniform distribution of sequencing errors.
@@ 4.3 Annotating the graph with copy numbers
For each member of the training set, we produce fake short-read sequencing data. For simplicity, this is single-ended (i.e. not in pairs from a larger template). They are randomly distributed from either strand with a random distribution of base-substitution errors; no indel errors are modelled.
Existing tools that take a short sequence, align it against a graph (GFA) and record the path include GraphAligner, vg giraffe and minigraph. We also produce our own kmer2node tool. All four tools annotate the GFA with edge-traversal counts and kmer counts within node sequences. The presence of repeats can still be problematic, as multiple paths may exist. It may not be possible to distinguish genuine sequence rearrangements from secondary hits arising from multiple copies of the repeated sequence in the pangenome. When reported, we filter out secondary alignments.
As mentioned above, our pangenome graph node sequences are a single representative or consensus sequence. However, the training genomes used to construct that pangenome may have many point mutations and small indels, which reduce kmer match rates. Kmer2node may use a list of recognition sequences, so we align the full-length training genomes to the pangenome using GraphAligner to produce a list of recognition sequences per node instead. This improves our kmer hit rate, but may also decrease the percentage that uniquely map. While indexing the GFA, we also get an estimate of the percentage of kmers we expect to be unique. This provides us with a kmer hit-rate expectation during alignment of short-read data, allowing us to improve the assessment of node sequence depth.
All the tools produce a GFA with kmer counts per node and frequencies of edge traversals. For pathfinder, this is sufficient as it can convert these to depth-normalised counts. To aid consistency, we also use pathfinder’s sub-graph connection analysis and depth-normalised counts as input to the QUBO solvers.
@@ 4.4 Classical post-processing: evaluating path solutions
Our path solvers produce a list of nodes and their orientation. From this, we can concatenate the sequences recorded in the GFA nodes, reverse-complementing as necessary, to produce a candidate genome sequence. Note that even if the path is perfect, this candidate genome may not be identical to the true genome, as our GFA files produced by minigraph collapse small variations into a single node. It is also possible that the true genome has novel structural rearrangements (copy-number changes or translocations) not present in the training data used to produce the pangenome graph, so a perfect path may not exist.
We align this candidate genome sequence against the known truth using bwa mem, reporting only primary and supplementary hits. This produces one or more alignments, each representing co-linear matches between the candidate and true genomes. The alignments are broadly equivalent to the contigs produced by a de novo assembler. Each alignment record may, in turn, have smaller variations (substitutions and small indels) that are not large enough to split the alignment into two. This provides a unified framework for comparing pangenome path reconstruction with more traditional de novo sequence assembly tools. We report the number of contigs and their N50 size. We align the contigs to the true genome to estimate the proportion of the true genome we have sampled, and vice versa. This two-way approach enables us to detect both over- and under-calling of copy-number variants.
These provide basic evidence on the contiguity of the assembled data, but not its accuracy. To this end, we also report the number of breaks when aligning these contigs back to the true genome (indicating incorrect path forks or large indels that require a supplementary alignment). These are analogous to false joins from a de novo assembler. For finer-grained evaluation, we also report small indels ($\geq$ 10 bp), regions of substantial base variation (multi-nucleotide polymorphism of 30% variance within a 100 bp window), and the overall percentage identity. Although BUSCO metrics are frequently used to assess assembly quality, these experiments assemble only small regions of synthetic genomes that do not contain any BUSCO genes, so it is not possible to use BUSCO for evaluation in this case.
We expect most users to prefer more contigs over incorrectly joined contigs, but the relative weighting of the two is a subjective decision that will depend on downstream usage.
A further optional refinement step is to realign the short reads to the generated candidate genome sequence and then take the consensus of the newly aligned sequences. This corrects for novel mutations in this sample or simplifications made during the minigraph pangenome creation step.
@@ 4.5 Formulating sequence assembly as an optimisation problem
A schematic representation of the process of converting a graph-traversal problem to a QUBO problem, with sample QUBO constraint terms highlighted.
@@
After annotation, our data is a graph with non-negative weights on the nodes and edges. Our current formulation only uses node weights, whereas tools such as pathfinder can use both sets of weights. Extending our methodology to include edge weights is a future direction that we discuss in section S2.
We formally define the pangenome graph annotated with copy numbers as a triplet $G=(V,E,w)$, where $V$ is the set of vertices, $E\subseteq V\times V$ is the set of edges and $\mathbbR+w:V\rightarrow\mathbb{R}_{+}$ is a weight function giving the estimated copy number of each vertex.
We then seek a walk on the graph that best explains the copy number data. This graph-traversal problem shares superficial similarities with the Travelling Salesman Problem. In particular, we define the following optimisation problem, which we refer to as Tangle Resolution due to the often complex, knot-like structures present in pangenomes. Note the implicit requirement in the problem that each step in $W$ traverses an edge in $G$.
@@
While this model serves as a useful starting point, it does not encompass all the biological features. Most notably, it fails to account for the double-stranded nature of genomic data. To better reflect the biology, we must consider a directed graph. The vertex set is now formed of 2 halves $V_{+},V_{-}$. The sequence data corresponding to $v_{-}\in V_{-}$ is the reverse complement of the sequence data corresponding to $v_{+}\in V_{+}$. Edges of the graph are directed and come in pairs: an edge $A+\rightarrow B+$ implies the existence of an edge $B-\rightarrow A$-, and vice versa. We call such a graph $G=(V_{+}\cup V_{-},E,w)$ oriented.
@@
Finally, for the setting of human data, we need to resolve diploid DNA. We therefore need to find a pair of walks $W_{1},\,W_{2}$ through an oriented graph that combine to explain the data. Formally, we have Diploid Tangle Resolution:
@@ 4.6 Mapping guided alignment to binary optimisation
[…]
QUBO is equivalent, under a linear change of variables, to the Ising model in statistical physics, a formulation extensively used in quantum annealing devices such as those developed by D-Wave Systems. As such, QUBO is widely employed in quantum computing contexts, where its unconstrained nature and binary domain align with the hardware capabilities of current quantum and neuromorphic processors. The primary advantages of QUBO include its flexibility in encoding diverse problems within a unified mathematical framework and its compatibility with both quantum and classical heuristic solvers. Its unconstrained nature simplifies solver implementation by eliminating the need for explicit constraint handling, instead requiring constraint satisfaction to be incorporated into the objective function through penalty terms.
Notably, however, the QUBO forms of many classic graph-traversal problems, such as the Travelling Salesman Problem, require quadratically many variables in the graph size, with each variable encoding whether the path visits a certain node at a certain time. QUBO formulations of these tasks traditionally use binary variables $x$ with 2 indices $t$ and $v$, with the variable $x_{t,v}=1$ if the solution path visits vertex $v$ at time $t$. The cost function is easily formulated by summing the cost of travelling between any 2 vertices $v$ and $v^{\prime}$, $c_{vv^{\prime}}$, over the edges that are traversed. This leads to an objective term $\sum_{t,v,v^{\prime}}c_{vv^{\prime}}\cdot x_{t,v}x_{t+1,v^{\prime}}$. For an assignment to represent a valid path, it is necessary to impose the constraint that exactly one vertex is visited at each time. This is accomplished through adding Lagrange multiplier terms to the cost function for each $t$, taking the form $\bigl(\sum_{v\in V}x_{t,v}-1\bigr)^{2}$.
@@
The total cost function consists of 3 parts: the first is a Lagrange multiplier term to enforce that the variable assignment encodes a walk that visits exactly 1 vertex at each time; the second is a multiplier term to enforce that, at each step, the walk traverses a graph edge, including travelling to or remaining in the virtual “end” node; the final term encodes the cost function $C_{G}(W)$.
@@
$C^{2}_{G}(\{x_{t,v}\})$ $=\Lambda_{2}\sum_{t=1}^{T-1}\left(\sum_{v,v^{\prime}inV}x_{t,v}x_{t+1,v^{\prime}}\mathbbm{1}_{\{(v,v^{\prime})\notin E\}}+\sum_{v\in V}x_{t,\text{end}}x_{t+1,v}\right),$ (5)
@@
There is some art to choosing the values of the Lagrange multipliers $\Lambda_{1},\Lambda_{2}$. If they are sufficiently large, then we can guarantee that the assignment that minimises $C_{G}(\{x_{t,v}\})$ encodes a walk that minimises $C_{G}(W)$. However, if they are too large, then the energy landscape will have steep valleys around assignments that satisfy the constraints, and changes in the cost function will be nearly indistinguishable to the optimiser. We find that intermediate-size values of $\Lambda_{1}=10$ and $\Lambda_{2}=5$ work well in practice. Empirically, for instances where copy number estimation was accurate, the penalties are small enough that the $C^{3}_{G}$ part is significant enough for the optimiser to find the true solution. Conversely, when copy number estimation is especially noisy, the penalties force a solution that obeys the constraints imposed by $C^{1}_{G}$ and $C^{2}_{G}$, overriding the inaccurate constraints of $C^{3}_{G}$.
To extend this method to Oriented Tangle Resolution, we simply introduce a pair of variables for each vertex and each time, resulting in $(2N+1)T$ variables. For Diploid Tangle Resolution, we further introduce a new set of variables $y_{t,v}$ for the second path, resulting in $2(2N+1)T$ variables. These variables have independent copies of $C^{1}_{G}$ and $C^{2}_{G}$ to enforce the constraints, but share a combined form of $C^{3}_{G}$.